# Relative Rényi Entropy Under Local Quenches in 2D CFTs

### v1

@@
Relative Rényi Entropy Under Local Quenches in 2D CFTs
Zi-Xuan Zhaoa,111zzx23@mails.jlu.edu.cn; The unusual ordering of authors instead of the standard alphabetical one in the hep-th community is to ensure that the student receives proper recognition for the student’s contribution under the current practice in Jilin University., Song Heb,c,d222hesong@nbu.edu.cn, Hao Ouyanga,333haoouyang@jlu.edu.cn, Hong-an Zenga,444zengha20@mails.jlu.edu.cn, Yu-Xuan Zhange,555zhangyuxuan@ucas.ac.cn
@@
bInstitute of Fundamental Physics and Quantum Technology, Ningbo University, Ningbo, Zhejiang 315211, China
cSchool of Physical Science and Technology, Ningbo University, Ningbo, 315211, China
@@
We study the relative Rényi entropy (RRE) under local quenches in two-dimensional conformal field theories (CFTs), focusing on rational CFTs (RCFTs) and holographic CFTs. In RCFTs, the RRE evolves as a monotonic function over time, depending on finite-dimensional matrices. It is sometimes symmetric, prompting an exploration of its relation to the trace squared distance. We also observe that relative entropy can fail to distinguish between operators, as it only captures information entering/exiting the subsystem. In holographic CFTs, an analytic continuation of the RRE reveals insights into the entanglement wedge, offering a new perspective on bulk geometry in AdS/CFT. Our results deepen the understanding of quantum information measures in RCFTs and holographic CFTs, highlighting connections to distinguishability and bulk reconstruction.
@@ 1 Introduction
Most previous studies have focused on the entanglement of a subsystem within a given quantum state [27, 28, 29, 30]. A natural extension is exploring whether other quantum information concepts can offer further insights when comparing two quantum states defined on the same Hilbert space. In this context, a particularly interesting quantity is the so-called relative entropy, which, for two given (reduced) density matrices $\rho$ and $\sigma$, is defined as:
$S(\rho||\sigma)=\text{tr}(\rho\log\rho)-\text{tr}(\rho\log\sigma).$ (1)
@@
In QFT, relative entropy is derived by modifying the replica trick for entanglement entropy [5], as introduced and refined by Lashkari [36, 37]. This method involves defining $\mathrm{Tr}(\rho\sigma^{k-1})$, which for integer $k$ serves as a generalized partition function or correlation function on an $k$-sheeted Riemann surface, thereby breaking the $Z_{k}$ symmetry among replicas. The relative entropy is then obtained through the replica limit $k\to 1$ of the so-called RRE defined as
$S_{k}(\rho||\sigma)=\frac{1}{k-1}\log\left[\text{tr}\rho^{k}-\text{tr}\rho\sigma^{k-1}\right],$ (2)
@@
provided that the parameter $k$ can be analytically continued from integer to complex values. This general method allows, at least in principle, the computation of relative entropy in any quantum field theory. However, to date, only a few direct calculations of relative entropy have been performed in 1+1-dimensional CFTs [36, 37, 39, 38, 42].
@@
This paper will investigate the RRE in RCFTs and holographic CFTs under local quenches. Although the information-theoretic properties of conformal descendants have been explored previously [50, 51, 52, 45], [28, 30] found that descendant operators and primary operators encode the same information in the context of entanglement/pseudo entropy, typically characterized by the quantum dimension $d_{\mathcal{O}}$ of the corresponding primary operator $\mathcal{O}$. Therefore, investigating the RRE under different local (descendant) quenches in RCFTs may deepen our understanding of the role of quantum information in CFTs by evaluating how distinguishable different quantum states are. Additionally, examining the information properties of conformal descendant operators in RCFTs through RRE may reveal their unique quantum information characteristics, aiding in distinguishing and understanding the information carried by different operators within the same conformal family. RRE may closely relate to geometric structures in bulk in the context of holographic duality since its special case, fidelity, can reconstruct the entanglement wedge [53]. Investigating RRE under local quenches may allow us to explore how quantum information measures can reconstruct bulk geometry, offering new perspectives and tools for the AdS/CFT correspondence.
@@ 2 Setup in 2D CFTs
$\rho=\frac{e^{-iHt}e^{-\epsilon H}|\psi\rangle\langle\psi|e^{\epsilon H}e^{iHt}}{\langle\psi|\psi\rangle},\quad\sigma=\frac{e^{-iHt^{\prime}}e^{-\epsilon H}|\psi^{\prime}\rangle\langle\psi^{\prime}|e^{\epsilon H}e^{iHt^{\prime}}}{\langle\psi^{\prime}|\psi^{\prime}\rangle},$ (3)
where $|\psi\rangle=\mathcal{O}(x)|\Omega\rangle$ and $|\psi^{\prime}\rangle=\mathcal{O}^{\prime}(x^{\prime})|\Omega\rangle$. Here, we introduce a small parameter $\epsilon$ to suppress high-energy modes [54].
By tracing out the degrees of freedom of $A^{c}$ (the complement of of subsystem $A$), we can obtain two reduced density matrices $\rho_{A}(t)=\text{tr}_{A^{c}}[\rho(t)]$ and $\sigma_{A}(t)=\text{tr}_{A^{c}}[\sigma(t)]$. The $k^{\rm th}$ RRE of $\rho_{A}$ and $\sigma_{A}$ can be derived through the replica method
$S_{k}(\rho_{A}||\sigma_{A})=\frac{1}{k-1}$ $\left(\text{tr}\rho_{A}^{k}-\text{tr}\rho\sigma_{A}^{k-1}\right)$

### v2

# Changchun 130012, People’s Republic of China

@@
Relative Rényi Entropy Under Local Quenches in 2D CFTs Zi-Xuan Zhaoa , Song Heb,c,d111hesong@nbu.edu.cn; Corresponding author, Hao Ouyanga,222haoouyang@jlu.edu.cn; Corresponding author, Hong-an Zenga,333zengha20@mails.jlu.edu.cn; Corresponding author, Yu-Xuan Zhange,444zhangyuxuan@ucas.ac.cn; Corresponding author
@@
bInstitute of Fundamental Physics and Quantum Technology, Ningbo University,
Ningbo, 315211, People’s Republic of China
cSchool of Physical Science and Technology, Ningbo University,
Ningbo, 315211, People’s Republic of China
@@
We study the relative Rényi entropy (RRE) under local quenches in two-dimensional conformal field theories (CFTs), focusing on rational CFTs (RCFTs) and holographic CFTs. In RCFTs, the RRE evolves as a monotonic function over time, depending on finite-dimensional matrices. It is sometimes symmetric, prompting an investigation into its relationship with the squared trace distance. We also observe that relative entropy can fail to distinguish between operators, as it only captures information entering/exiting the subsystem. In holographic CFTs, an analytic continuation of the RRE reveals insights into the entanglement wedge, offering a new perspective on bulk geometry in AdS/CFT. Our results deepen the understanding of quantum information measures in RCFTs and holographic CFTs, highlighting connections to distinguishability and bulk reconstruction.
@@ 1 Introduction
Most previous studies have focused on the entanglement of a subsystem within a given quantum state, for instance, the dynamics of subsystem entanglement entropy for a non-equilibrium state created by local quenches [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56] or global quenches [57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80] protocols. A natural extension is exploring whether other quantum information concepts can offer further insights when comparing two quantum states defined on the same Hilbert space. In this context, a particularly interesting quantity is the so-called relative entropy, which, for two given (reduced) density matrices $\rho$ and $\sigma$, is defined as:
$S(\rho||\sigma)=\text{tr}[\rho\log\rho]-\text{tr}[\rho\log\sigma].$ (1)
@@
In QFT, relative entropy is derived by modifying the replica trick for entanglement entropy [5], as introduced by Lashkari [86]. This method involves defining the so-called sandwiched relative Rényi entropy [100, 101]
$S_{\alpha}(\rho||\sigma)=\begin{cases}\frac{1}{\alpha-1}\log\text{tr}([\sigma^{\frac{1-\alpha}{2\alpha}}\rho\sigma^{\frac{1-\alpha}{2\alpha}}]^{\alpha})\quad\text{if}\rho\not\perp\sigma\wedge(\sigma\gg\rho\vee\alpha<1)\\ \infty\quad\text{else}\end{cases}$ (2)
for any $\alpha\in(1/2,1)\cup(1,\infty)$. For the range $\alpha\in(0,1/2)$, it is more natural to use the Petz $\alpha$-relative entropy [102] defined as
$S_{\alpha}(\rho||\sigma)=\frac{1}{\alpha-1}\log\text{tr}[\rho^{\alpha}\sigma^{1-\alpha}].$ (3)
@@
An alternative approach [87] used for computing the matrix elements of the modular Hamiltonian corresponding to excited states involves introducing a Rényi-like one-parameter quantity $\mathrm{Tr}(\rho\sigma^{k-1})$, which for integer $k$ serves as a generalized partition function or correlation function on an $k$-sheeted Riemann surface, thereby breaking the $Z_{k}$ symmetry among replicas. The relative entropy is then obtained through the replica limit $k\to 1$ of the RRE [^5] For convenience, we refer to the following Rényi-like one-parameter quantity as RRE throughout this work. defined as
$S_{k}(\rho||\sigma)=\frac{1}{k-1}\left(\log\text{tr}[\rho^{k}]-\log\text{tr}[\rho\sigma^{k-1}]\right),$ (4)
@@
provided that the parameter $k$ can be analytically continued from integer to complex values. This general method, at least in principle, allows the computation of relative entropy in any quantum field theory. However, to date, only a few direct calculations of relative entropy have been performed in 1+1-dimensional CFTs [86, 87, 89, 88, 92].
While the RRE does not require certain formal properties of the sandwiched RRE—such as UV finiteness, non-negativity, and satisfaction of the data processing inequality [100] [^6] We sincerely thank the anonymous referee for highlighting this important point.— and it is still unknown whether this quantity has a quantum information interpretation, it remains worthy of investigation because it provides a more intuitive realization of the replica method while circumventing the subtle analytical continuation required when extending from positive integer to fractional values. Our preference for RRE in this work does not imply its superiority over sandwiched RRE or Petz $\alpha$-RRE. The RRE was chosen purely for its clearer replica geometry visualization and more straightforward computational implementation in our framework. While one may alternatively adopt sandwiched RRE or Petz $\alpha$-RRE formulations, our computational framework demonstrates they would yield equivalent results.
@@
This paper will investigate the RRE in RCFTs and holographic CFTs under local quenches. Although the information-theoretic properties of conformal descendants have been explored previously [103, 36, 104, 95, 37, 105], it has been found that operators in the same conformal family encode the same information in the context of entanglement entropy [36, 37] or pseudo entropy [106, 105], typically characterized by the quantum dimension $d_{\mathcal{O}}$ of the corresponding primary operator $\mathcal{O}$. In [36], the authors pointed out that when considering the Schmidt decomposition of quantum states, different descendant states may share the same Schmidt coefficients while possessing different Schmidt bases, leading to identical entanglement entropy. This information, arising from the change in the Schmidt basis, can be captured by the RRE. Therefore, investigating the RRE under different local (descendant) quenches in RCFTs may deepen our understanding of the role of quantum information in CFTs by evaluating how distinguishable different quantum states are. Additionally, examining the information properties of conformal descendant operators in RCFTs through RRE may reveal their unique quantum information characteristics, aiding in distinguishing and understanding the information carried by different operators within the same conformal family. RRE may closely relate to geometric structures in bulk in the context of holographic duality since its special case, fidelity, can reconstruct the entanglement wedge [107]. Investigating RRE under local quenches may allow us to explore how quantum information measures can reconstruct bulk geometry, offering new perspectives and tools for the AdS/CFT correspondence.
@@ 2 Setup in 2D CFTs
$\rho=\frac{e^{-iHt}e^{-\epsilon H}|\psi\rangle\langle\psi|e^{-\epsilon H}e^{iHt}}{\langle\psi|e^{-2\epsilon H}|\psi\rangle},\quad\sigma=\frac{e^{-iHt^{\prime}}e^{-\epsilon H}|\psi^{\prime}\rangle\langle\psi^{\prime}|e^{-\epsilon H}e^{iHt^{\prime}}}{\langle\psi^{\prime}|e^{-2\epsilon H}|\psi^{\prime}\rangle},$ (5)
where $|\psi\rangle=\mathcal{O}(x)|\Omega\rangle$ and $|\psi^{\prime}\rangle=\mathcal{O}^{\prime}(x^{\prime})|\Omega\rangle$. Note that an infinitesimal positive parameter $\epsilon$ has been introduced to suppress high-energy modes [57].
By tracing out the degrees of freedom of $A^{c}$ (the complement of the subsystem $A$), we can obtain two reduced density matrices $\rho_{A}(t)=\text{tr}_{A^{c}}[\rho(t)]$ and $\sigma_{A}(t)=\text{tr}_{A^{c}}[\sigma(t)]$. The $k^{\rm th}$ RRE of $\rho_{A}$ and $\sigma_{A}$ can be derived through the replica method
$S_{k}(\rho_{A}||\sigma_{A})=\frac{1}{k-1}$ $\left(\log\text{tr}[\rho_{A}^{k}]-\log\text{tr}[\rho\sigma_{A}^{k-1}]\right)$